# GateRA: Token-Aware Modulation for Parameter-Efficient Fine-Tuning

### v1

@@ Introduction
Large pre-trained models have revolutionized natural language processing and vision-language tasks, yet their massive parameter counts make full fine-tuning increasingly impractical. Parameter-efficient fine-tuning (PEFT) methods, such as LoRA (Hu et al. 2021a), DoRA (Liu et al. 2024), and HiRA (Huang et al. 2025), address this by injecting trainable low-rank updates into frozen backbones, enabling task adaptation with minimal parameter overhead.
@@
(a) Query projection layer in self-attention.
(b) Key projection layer in self-attention.
(c) Up-projection layer in the MLP block.
@@
Parameter-efficient fine-tuning (PEFT) methods aim to adapt large-scale pre-trained models to downstream tasks while minimizing the number of trainable parameters. Early approaches include adapter tuning (Houlsby et al. 2019), prefix-tuning (Li and Liang 2021), and prompt-tuning (Lester, Al-Rfou, and Constant 2021), which insert task-specific modules or learn prompt embeddings. LoRA (Hu et al. 2021a) further improves efficiency by injecting trainable low-rank matrices into weight projections, achieving strong performance across a range of language and vision tasks. While these methods reduce memory and compute costs, they typically apply uniform updates across all tokens, lacking the flexibility to differentiate token-level importance. In contrast, our method introduces a token-aware gating mechanism that modulates the update strength dynamically per-token, allowing fine-grained control over adaptation.
@@
Numerous LoRA variants have been proposed to improve flexibility and robustness. AdaLoRA (Zhang et al. 2023) adapts rank allocation across layers based on sensitivity; QLoRA (Dettmers et al. 2023) combines LoRA with quantization for low-resource fine-tuning. DoRA (Liu et al. 2024) decomposes weights into direction and magnitude for improved gradient flow, while MoRA (Zhao et al. 2023) applies a multiplicative gating mechanism. LoRA-Slow (Zhang et al. 2024) introduces gradual low-rank updates, and other efforts explore structure-aware improvements such as tensorized (Li et al. 2023), Fourier-based (Gao et al. 2024), or expert-based (Chen et al. 2024) designs. Our method is orthogonal and complementary to these extensions, focusing on dynamic token-wise modulation that can be applied on top of additive (LoRA), directional (DoRA), or multiplicative (HiRA) PEFT variants.
@@
A few recent works have begun to explore input-dependent or sparse adaptation strategies. AdaMix (Xiong et al. 2025) uses instance-wise adapter selection via a routing mechanism; UNIPELT (Mao et al. 2021) learns a task-specific gating of multiple adaptation modules. In vision, SPDF (Thangarasa et al. 2023) selectively activates PEFT modules based on token importance. These methods introduce task-level or layer-level control, but rarely operate at the fine-grained token level during inference. Our approach differs by directly modeling token-level update decisions through a gating function trained end-to-end, and further regularized to produce sparse, interpretable modulation across time.
@@
Parameter-efficient fine-tuning (PEFT) enables the adaptation of large pre-trained models by injecting small, trainable modules into frozen backbones. Among various PEFT strategies, HiRA (Huang et al. 2025) has demonstrated strong performance by applying multiplicative low-rank updates to the weight matrix. Specifically, HiRA modifies the pre-trained weight $W_{0}$ as follows:
@@
While the gating mechanism introduced in GateRA enables token-level modulation of adaptation strength, unconstrained learning of $g(x)$ may lead to overly smooth or indecisive gating behaviors. In practice, we observe that without additional regularization, the model tends to assign ambiguous gating values (e.g., close to 0.5), resulting in uniformly soft updates across all tokens and diminishing the benefits of selective adaptation.
@@
Table 1: Accuracy comparison on commonsense reasoning tasks. We report results on eight subtasks from the CommonsenseQA benchmark. GateRA consistently outperforms all baselines under both LLaMA-2-7B and LLaMA-3-8B.
@@
We follow (Hu et al. 2021a) and adopt eight widely used sub-tasks with predefined training and test splits, totaling 170,420 query-answer pairs. The sub-tasks include: BoolQ(Clark et al. 2019): yes/no questions; PIQA(Bisk et al. 2020): physical commonsense; SIQA(Sap et al. 2019): social reasoning; HellaSwag(Zellers et al. 2019): sentence completion; WinoGrande(Sakaguchi et al. 2021): coreference-based commonsense; ARC-e and ARC-c(Clark et al. 2018): multiple-choice science QA; OBQA (Mihaylov et al. 2018): multi-step reasoning questions. These datasets collectively cover binary, multiple-choice, and cloze-style questions, posing different levels of difficulty for LLMs.
@@
We use the ConvAI2 dataset (Dinan et al. 2019), which consists of 17,878 multi-turn dialogues for training and 1,000 for testing. Each instance includes persona profiles and a conversational history. Following (Liu et al. 2020; Song et al. 2021; Huang et al. 2023), we adopt a self-persona setting where only the speaker’s persona is visible.
@@
We evaluate symbolic and multi-step reasoning capabilities on GSM8K (Cobbe et al. 2021), a challenging benchmark consisting of grade-school math word problems requiring structured solution steps.
@@
For commonsense reasoning, we follow the keyword matching protocol in (Hu et al. 2021a; Liu et al. 2024), using accuracy as the primary metric. The decoded answer is scanned for specific keywords (e.g., “true”, “B”), and the first match is taken as the prediction. If no match is found, the answer is considered incorrect. This rule-based scheme enables consistent evaluation across different question formats. For ConvAI2, we report BLEU (Papineni et al. 2002) and BERTScore (Zhang et al. 2019) to measure generation quality and semantic similarity. For GSM8K, we report accuracy by comparing the model’s final numeric prediction with the ground truth using an exact match.
@@
We compare GateRA against strong PEFT methods including LoRA (Hu et al. 2021b), DoRA (Liu et al. 2024), MoRA (Jiang et al. 2024), and HiRA (Huang et al. 2025). To ensure fair comparison, we use the same injection locations (query/key/value/MLP) and equal rank $r=16$. For prompt-based baselines like Prompt Tuning (Lester, Al-Rfou, and Constant 2021) and P-Tuning (Liu et al. 2021), we adjust the soft prompt length to maintain comparable trainable parameter budgets.
@@
Table 2: Results on the CONVAI2 dialogue generation task. We evaluate BLEU, BERT-based F1/Recall/Precision, Meteor, and ROUGE-L. GateRA achieves the best performance under both LLaMA-2-7B and LLaMA-3-8B, outperforming all prompt and PEFT baselines.
@@
Following (Liu et al. 2024), we evaluate on LLaMA-2-7B (Touvron et al. 2023) and LLaMA-3-8B (Grattafiori et al. 2024). All models are fine-tuned using the AdamW optimizer (Loshchilov and Hutter 2017) with a learning rate of 1e-3 and linear warmup of 100 steps. Each model is trained for 3 epochs (2 epochs for math reasoning), and the best checkpoint is selected based on validation performance. For each run, we report average results over 5 random seeds.
@@
We evaluate the performance of GateRA on mathematical reasoning tasks using the MetaMath dataset (Yu et al. 2023) for training and GSM8K for evaluation. As reported in Table 3, GateRA achieves state-of-the-art performance compared to all PEFT baselines under the LLaMA-3-8B setting. Specifically, GateRA attains an accuracy of 72.11%, outperforming HiRA (70.81%), MoRA (67.98%), DoRA (66.12%), and LoRA (65.89%). These findings demonstrate that GateRA is particularly effective for complex symbolic reasoning, such as multi-step mathematical tasks, by leveraging token-aware modulation to enhance expressiveness while maintaining parameter efficiency.
@@
Table 3: Results on mathematical reasoning tasks. We evaluate models on the GSM8K benchmark with MetaMath (Yu et al. 2023) training. GateRA achieves the best performance under LLaMA-3-8B, outperforming all PEFT baselines.
@@
Figure 3: Token-adaptive performance under different ranks. We compare GateRA’s performance across eight commonsense reasoning datasets using two configurations: rank=1616 and rank=3232.
@@ Conclusion
We propose GateRA, a unified and lightweight framework for test-time adaptation in PEFT methods. GateRA introduces a token-aware modulation mechanism that dynamically adjusts the strength of prompt updates based on epistemic uncertainty, enabling selective and input-specific adaptation. This allows the model to concentrate updates on uncertain or task-critical tokens while preserving generalizable representations. We further connect the gating function to a spike-and-slab prior and introduce entropy-based regularization to promote sparse and confident adaptation. Theoretically, GateRA induces a soft masking effect on gradient flow, improving the balance between stability and adaptability. Extensive experiments across multiple NLP benchmarks demonstrate that GateRA consistently improves over PEFT baselines.

### v2

@@ Introduction
Large pre-trained models have revolutionized natural language processing and vision-language tasks, yet their massive parameter counts make full fine-tuning increasingly impractical. Parameter-efficient fine-tuning (PEFT) methods, such as LoRA (lora_original), DoRA (liu2024dora), and HiRA (huang2025hira), address this by injecting trainable low-rank updates into frozen backbones, enabling task adaptation with minimal parameter overhead.
@@
Parameter-efficient fine-tuning (PEFT) methods aim to adapt large-scale pre-trained models to downstream tasks while minimizing the number of trainable parameters. Early approaches include adapter tuning (houlsby2019parameter), prefix-tuning (li2021prefix), and prompt-tuning (lester2021power), which insert task-specific modules or learn prompt embeddings. LoRA (lora_original) further improves efficiency by injecting trainable low-rank matrices into weight projections, achieving strong performance across a range of language and vision tasks. While these methods reduce memory and compute costs, they typically apply uniform updates across all tokens, lacking the flexibility to differentiate token-level importance. In contrast, our method introduces a token-aware gating mechanism that modulates the update strength dynamically per-token, allowing fine-grained control over adaptation.
@@
Numerous LoRA variants have been proposed to improve flexibility and robustness. AdaLoRA (zhang2023adaptive) adapts rank allocation across layers based on sensitivity; QLoRA (dettmers2023qlora) combines LoRA with quantization for low-resource fine-tuning. DoRA (liu2024dora) decomposes weights into direction and magnitude for improved gradient flow, while MoRA (jiang2024mora) employs a square matrix for high-rank adaptation. LoRASC (li2024lorasc) utilizes cascaded learning and slow-fast updates, and other efforts explore structure-aware improvements such as tensorized (yang-etal-2024-loretta), Fourier-based (borse2024foura), or expert-based (zhang2024milora) designs. Our method is orthogonal and complementary to these extensions, focusing on dynamic token-wise modulation that can be applied on top of additive (LoRA), directional (DoRA), or multiplicative (HiRA) PEFT variants.
@@
A few recent works have begun to explore input-dependent or sparse adaptation strategies. Apart (qi2025adaptive) uses instance-wise adapter selection via a routing mechanism; UNIPELT (mao2022unipelt) learns a task-specific gating of multiple adaptation modules. In vision, TR-PTS (luo2025tr) selectively activates PEFT modules based on token and task. These methods introduce task-level or layer-level (yao2024layer) control, but rarely operate at the fine-grained token level during inference. Our approach differs by directly modeling token-level update decisions through a gating function trained end-to-end, and further regularized to produce sparse, interpretable modulation across time.
@@
Parameter-efficient fine-tuning (PEFT) enables the adaptation of large pre-trained models by injecting small, trainable modules into frozen backbones. Among various PEFT strategies, HiRA (huang2025hira) has demonstrated strong performance by applying multiplicative low-rank updates to the weight matrix. Specifically, HiRA modifies the pre-trained weight $W_{0}$ as follows:
@@
While the gating mechanism introduced in GateRA enables token-level modulation of adaptation strength, unconstrained learning of $g(x)$ may lead to overly smooth or indecisive gating behaviors. In practice, we observe that without additional regularization, the model tends to assign ambiguous gating values (e.g., close to 0.5) to some tokens, resulting in uniformly soft updates across tokens and diminishing the benefits of selective adaptation.
@@
Table 1: Accuracy comparison on commonsense reasoning tasks. We report results on eight subtasks from the CommonsenseQA benchmark. GateRA consistently outperforms all baselines under both LLaMA-2-7B (L2-7B) and LLaMA-3-8B (L3-8B).
@@
We follow (hu2023llm) and adopt eight widely used sub-tasks with predefined training and test splits, totaling 170,420 query-answer pairs. The sub-tasks include: BoolQ (clark2019boolq): yes/no questions; PIQA (bisk2020piqa): physical commonsense; SIQA (sap2019socialiqa): social reasoning; HellaSwag (zellers2019hellaswag): sentence completion; WinoGrande (sakaguchi2021winogrande): coreference-based commonsense; ARC-e and ARC-c(clark2018think): multiple-choice science QA; OBQA (mihaylov2018can): multi-step reasoning questions. These datasets collectively cover binary, multiple-choice, and cloze-style questions, posing different levels of difficulty for LLMs.
@@
We use the ConvAI2 dataset (dinan2019second), which consists of 17,878 multi-turn dialogues for training and 1,000 for testing. Each instance includes persona profiles and a conversational history. Following (liu2020you; song2021bob; huang2023personalized), we choose a self-persona configuration where only the speaker’s persona is visible.
@@
We evaluate symbolic and multi-step reasoning capabilities on GSM8K (cobbe2021training), a challenging benchmark consisting of grade-school math word problems requiring structured solution steps.
@@
For commonsense reasoning, we follow the keyword matching protocol in (liu2024dora; huang2025hira), using accuracy as the primary metric. The decoded answer is scanned for specific keywords (e.g., “true”, “B”), and the first match is taken as the prediction. If no match is found, the answer is considered incorrect. This rule-based scheme enables consistent evaluation across different question formats. For ConvAI2, we report BLEU (papineni2002bleu) and BERTScore (zhang2019bertscore) to measure generation quality and semantic similarity. For GSM8K, we report accuracy by comparing the model’s final numeric prediction with the ground truth using an exact match.
@@
We compare GateRA against strong PEFT methods including LoRA (hu2021lora), DoRA (liu2024dora), MoRA (jiang2024mora), and HiRA (huang2025hira). To ensure fair comparison, we use the same injection locations (query/key/value/MLP) as HiRA.
@@
Table 2: Results on the CONVAI2 dialogue generation task. We evaluate BLEU, BERT-based F1/Recall/Precision, Meteor, and ROUGE-L.
@@
Following (liu2024dora), we evaluate on LLaMA-2-7B (touvron2023llama) and LLaMA-3-8B (grattafiori2024llama). All models are fine-tuned using the AdamW optimizer (loshchilov2017decoupled) with a learning rate of 1e-3 and warmed up with 100 steps.
@@
We evaluate the performance of GateRA on mathematical reasoning tasks using the MetaMath dataset  (yu2023metamath) for training and GSM8K for evaluation. As reported in Table 3, GateRA achieves state-of-the-art performance compared to all PEFT baselines under the LLaMA-3-8B setting. Specifically, GateRA attains an accuracy of 72.11%, outperforming HiRA (70.81%), MoRA (67.98%), DoRA (66.12%), and LoRA (65.89%). These findings demonstrate that GateRA is effective for complex reasoning, by leveraging token-aware modulation to enhance expressiveness while maintaining parameter efficiency.
@@
Table 3: Results on mathematical reasoning tasks. We evaluate models on the GSM8K benchmark with MetaMath (yu2023metamath) training.
@@
Figure 3: Token-adaptive performance under different ranks.
@@ Conclusion
We propose GateRA, a unified and lightweight framework for test-time adaptation in PEFT methods. GateRA introduces a token-aware modulation mechanism that dynamically adjusts the strength of magnitude of weight usage based on epistemic uncertainty, enabling selective and input-specific adaptation. This allows the model to concentrate updates on uncertain or task-critical tokens while preserving generalizable representations. We further connect the gating function to a spike-and-slab prior and introduce entropy-based regularization to promote sparse and confident adaptation. Theoretically, GateRA induces a soft masking effect on gradient flow, improving the balance between stability and adaptability. Extensive experiments across multiple NLP benchmarks demonstrate that GateRA consistently improves over PEFT baselines.